OpenStudy (mackenzie2013):

Yehudy invested $19,250 in a simple interest account. After 6 years and no additional deposits or withdrawals, the account balance was $22,957.55. Find the interest rate for the account. Round to the nearest hundredth of a percent if needed. 32.1% 1.99% 19.88% 3.21%

OpenStudy (justnick09):

19.88

OpenStudy (justnick09):

do u want me to tell u how i got it

OpenStudy (mackenzie2013):

sure

OpenStudy (justnick09):

ok so its 6 years right?

OpenStudy (mackenzie2013):

Yes

OpenStudy (justnick09):

\[22957.55/6\]

OpenStudy (justnick09):

that gives u 3826.258

OpenStudy (justnick09):

now you divide that by the initial amount of 19250 nd boom

OpenStudy (mackenzie2013):

Oh, I see. Thank you!

OpenStudy (justnick09):

np

OpenStudy (johnweldon1993):

Are we sure there? So recall the formula for simple interest I = PRT I = interest earned P = principal amount R = rate of interest T = time in years So if we DO have 19.88% interest rate...that would give us... \[\large I = 19250\times .1988 \times 6 = 22961.4\] Now remember, this is just the interest earned! so adding that to the principal we would have 42211.4 in the account...would be nice, but sadly it is incorrect If we want to calculate the true interest rate...we use the formula \[\large A = P(1 + rt)\] Where A is the amount after all the interest is earned *which we are given* So plugging everything in we would have \[\large 22957.55 = 19250(1 + r(6))\] Now divide both sides by 19250 \[\large 1.1926 = 1 + r(6)\] Subtract 1 from both sides \[\large 0.1926 = r(6)\] and finally divide both sides by 6 \(\large r = 0.0321\) or in percent form \(\large 3.21\text{%}\)